# Detecting Functional Bugs in Smart Contracts through LLM-Powered and Bug-Oriented Composite Analysis

### v1

@@ 4.3. Invariant Checker Generation Accuracy
As shown in Table 5, the invariant checker generation module demonstrates excellent performance in critical variable and principal statement extraction, achieving an overall precision of $98.18\%$. Specifically, for the variable extraction task, PromFuzz attains a perfect precision of $100\%$. For the statement extraction task, it achieves a precision of $83.33\%$, with only one false positive. This false positive arises from the complex business behaviors in smart contracts, where PromFuzz fails to differentiate between multiple similar behaviors in a single complex case. These critical variables and principal statements are expected to derive $23$ correct invariant checkers but ultimately generate $22$ correct ones, resulting in a precision of $95.65\%$.
Table 5. Variables and statements extraction accuracy.
@@
Variables Extraction 49 0 100%
Statements Extraction 5 1 83.3%
Overall 54 1 98.18%
Table 6. Functional bug detection accuracy.
Tool $D_{Exploit-SC}$ $D_{BV-SC}$ $D_{Synth-SC}$ Overall
TP FP FN TP FP FN TP FP FN TP FP FN Recall F1-Score
PromFuzz 6 0 1 8 0 2 6 0 0 20 0 3 86.96% 93.02%
GPTScan 2 4 5 5 4 5 0 4 6 7 12 16 30.43% 33.33%
SMARTINV 4 19 3 1 20 9 1 4 5 6 43 17 26.09% 16.67%
ItyFuzz 0 0 7 0 0 10 0 0 6 0 0 23 0 0
SMARTIAN 0 0 6 1 0 8 2 0 2 3 0 16 15.79% 27.27%
4.4. Functional Bug Detection Accuracy
@@
This subsection answers RQ3. In this step, we evaluate the functional bug detection accuracy of PromFuzz with two metrics: recall and F1-score. We compare PromFuzz with four off-the-shelf tools: GPTScan, SMARTINV, ItyFuzz, and SMARTIAN (Choi et al., 2021). We do not compare PromFuzz with several program verification-based techniques (Kalra et al., 2018; Permenev et al., 2020; Duan et al., 2022), since they are largely ineffective in detecting functional bugs, as their constraints are designed for implementation bugs. Both ItyFuzz and SMARTIAN are fuzzers designed for detecting bugs in smart contracts. We perform them on $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$, respectively.
@@
As shown in Table 6, PromFuzz demonstrates superior performance in detecting functional bugs across three datasets. On $D_{Exploit-SC}$, PromFuzz identifies 6 true positives with no false positives and only 1 false negative. Similarly, on $D_{BV-SC}$, it identifies 8 true positives, with no false positives and 2 false negatives. For $D_{Synth-SC}$, PromFuzz perfectly identifies 6 true positives with no false negatives or false positives. Cumulatively, PromFuzz achieves an overall recall of 86.96% and an F1-score of 93.02%, detecting 20 true positives and maintaining 0 false positives across all datasets, with only 3 false negatives. Further analysis of the false negatives identified by PromFuzz reveals two primary causes. First, one false negative results from inaccuracies in the extraction of critical variables and principal statements. Second, the remaining two false negatives occur during the LLM-driven analysis step, where GPT fails to identify two critical bugs. Besides, with each smart contract project scan, PromFuzz takes an average of 43.8 seconds and costs approximately \$0.15, demonstrating its great efficiency and cost-effectiveness.
@@
In comparison, GPTScan and SMARTINV yield similar numbers of true positives and false negatives. However, SMARTINV records the highest number of false positives among the five tools assessed. Furthermore, ItyFuzz fails to identify functional bugs without the specific invariants we provide. SMARTIAN, another fuzzing tool, shows inconsistent performance as it fails to operate on several smart contract files. Consequently, we exclude SMARTIAN from testing against all bugs. We further analyze the false positives and false negatives of these tools. The reasons behind the false positives for both GPTScan and SMARTINV are already discussed in Section 4.2. The false negatives for GPTScan mainly arise because its prompts do not fully cover several specific functional bugs, such as unauthorized behavior, despite considerable efforts. SMARTINV experiences false negatives due to its lack of comprehensive reasoning capabilities, resulting in some bugs being overlooked. Both ItyFuzz and SMARTIAN are limited by their original oracles, which lack the capability to identify functional bugs effectively.
@@
Table 7. Zero-day functional bug detection results. TVL represents the total value of cryptocurrency assets locked in the respective DeFi projects.
@@ 5 Yield & Sericves \$ 3,700M 5 5
4.5. Zero-day Functional Bug Detection
This subsection answers RQ4. We perform PromFuzz on $D_{Real-DeFi}$ to evaluate its performance in detecting zero-day functional bugs in real-world smart contracts. As illustrated in Table 7, PromFuzz successfully identifies $30$ zero-day functional bugs across $6$ DeFi projects. We have reported all detected bugs to the respective vendors, and $24$ of these have been assigned CVE IDs. The total market value of DeFi projects affected by these bugs is estimated at \$$\$18.2$ billion. As these bugs have not yet been fully resolved, we maintain confidentiality by anonymizing the details of both the vendors and the specific CVE information. Moreover, we notice that zero-day bugs primarily fall into four categories: 13 bugs involve AMM price oracle manipulation, $7$ concern non-AMM price oracle manipulation, $4$ are related to wrong checkpoint order, $4$ pertain to risky first deposit, and $2$ involve wrong interest rate order. Further analysis in Section 5 delves into the practical impacts of these bugs, examining specific cases to illustrate how attackers exploit these bugs.
5. Case Study
To obtain an in-depth understanding of the practical impacts of functional bugs, we detail two such bugs identified by PromFuzz in DeFi projects.
@@ 5. Case Study
5 if (_amount > type(uint128).max || _amount == 0 || share == 0){
6 revert InvalidAmount();
7 }
@@
13 if (totalPooledEther == 0) {
14 return _depositAmount;
15 }
@@
17 }
18}
Figure 8. The Risky First Deposit bug (line 13-15).
Risky First Deposit. As shown in Figure 8, the code contains a risky first deposit bug that can lead to an unfair advantage for the first user and a distorted distribution of shares. Specifically, at line 14, when totalPooledEther == 0, the share variable at line 5 is set equal to _depositAmount, since the return value is assigned to _depositAmount in the _sharesForDepositAmount() function. Subsequent depositors have their shares calculated by (_depositAmount $\times$ eETH.totalShares()) / totalPooledEther, which depends on the values of eETH.totalShares() and totalPooledEther. While the calculation logic seems standard for compliant depositors, the first depositor can exploit the system by manipulating totalPooledEther. We can front-run other depositors’ transactions and inflate the price of pool tokens through a substantial “donation.” For instance, a malicious early user could deposit 1 wei of asset tokens as the first depositor, receiving 1 wei of shares tokens. Subsequently, the malicious sends 1 ETH to the pool, inflating totalPooledEther to 1 ETH + 1 wei (equal to ($11e18+$ 1) wei). In this scenario, the second depositor, depositing 2 ETH (equal to $182e$ 18 wei) of asset tokens, would receive only 1 wei of shares tokens. Because 1.99 wei of shares tokens (calculated by $2e18/(1e18+1)$) will be rounded down to 1 wei of shares tokens in smart contracts. They will lose 0.5 ETH if they withdraw right after the _deposit(). This disproportionate calculation of the first user’s shares can disadvantage subsequent users who receive fewer shares relative to their deposits.
AMM Price Oracle Manipulation. As shown in Figure 9, the code contains an AMM price oracle manipulation bug, which arises from the method of calculating token prices based solely on the supplies of two tokens. Specifically, _reserveA and _reserveB represent the respective reserve amounts of two assets. In line 15 of the code, the amount of token B is calculated using the amount of token A, along with the reserves of token A and token B. However, if the reserve amount of either token is maliciously manipulated, it can distort the price and lead to significant price fluctuations. For instance, consider token exchanges in the smart contract using the PoolHelpers library. Initially, the pool contains 500 token A and 500 token B. An attacker exchanges 100 token A for 100 token B, changing the pool to 600 token A and 400 token B. Then, the second user exchanges 30 token A and can only get 20 token B in return, leaving the pool with 630 token A and 380 token B. Finally, the attacker uses the 100 token B they obtained to exchange for 165 token A, ultimately making a profit of 65 token A.
@@
16 }
17}
Figure 9. The AMM Price Oracle Manipulation bug (line 15).
6. Discussion
@@ 7. Related work
Recent efforts to evaluate the security of smart contracts have produced numerous methodologies. These approaches are broadly classified into three categories: static analysis, dynamic analysis, and AI-based analysis.
@@
Static Analysis. Tools like Slither (Feist et al., 2019), Zeus (Kalra et al., 2018), Securify (Tsankov et al., 2018), Madmax (Grech et al., 2018), Invcon (Liu and Li, 2022), Verismart (So et al., 2020), and Verisol (Wang et al., 2019a) analyze smart contract code statically. Slither converts Solidity code into SlithIR to find bugs. Securify uses compliance and violation patterns to assess code safety. Madmax decompiles EVM bytecode and uses a logic-based specification for bug detection. Verisol translates Solidity into Boogie, formalizing semantic conformance against state machine workflows and reducing semantic checking to safety verification. However, these tools primarily focus on implementation bugs like integer overflow and reentrancy, and are less effective at identifying functional bugs.
@@
Dynamic Analysis. Tools like ItyFuzz (Shou et al., 2023), Smartian (Choi et al., 2021), Harvey (Wüstholz and Christakis, 2020), Echidna (Grieco et al., 2020), ContractFuzzer (Jiang et al., 2018), Mythril (myt, 2024), Sailfish (Bose et al., 2022), Manticore (Mossberg et al., 2019), Teether (Krupp and Rossow, 2018), and Oyente (Luu et al., 2016) analyzes programs during execution. ItyFuzz uses snapshot-based fuzzing to reduce re-execution overhead, prioritizing intriguing snapshot states. Smartian combines static analysis for seed generation with feedback mechanisms. Sailfish identifies state-inconsistency bugs using a combination of lightweight exploration and symbolic evaluation. While dynamic analysis captures more realistic behaviors by executing smart contracts, it still primarily targets implementation bugs and struggles to effectively detect functional bugs.
@@
AI-based Analysis. SmarTest (So et al., 2021), ILF (He et al., 2019),xFuzz (Xue et al., 2024), and GNN-based tools (Zhuang et al., 2020; Liu et al., 2021, 2023) use AI to enhance smart contract bug detection. SmarTest combines symbolic execution with a language model to prioritize bugs. GNN-based tools leverage graph neural networks with expert patterns. While these tools improve detection capabilities, they mainly address implementation bugs. Recently, LLM-based tools like GPTScan (Sun et al., 2024) and SMARTINV (Wang et al., 2024) have been used to address functional bugs by analyzing discrepancies between business logic and code. These works, including the PromFuzz introduced here, open new avenues for detecting functional bugs in smart contracts using AI’s cognitive capabilities.
8. Conclusion
@@
In this paper, we introduce PromFuzz, an automated and practical system designed to detect functional bugs in smart contracts, which bridges the gap between understanding the high-level business model logic and scrutinizing the low-level implementations in smart contracts. PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, representing a significant performance improvement over state-of-the-art methods. Additionally, we have created the first ground-truth dataset for functional bugs, including $23$ verified functional bugs from various DeFi projects. Furthermore, our in-depth analysis of $6$ real-world DeFi projects has identified $30$ zero-day bugs, with $24$ of them receiving CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.

### v2

@@ IV-C Invariant Checker Generation Accuracy
As shown in Table IV, the invariant checker generation module demonstrates excellent performance in critical variable and principal statement extraction, achieving an overall precision of $94.92\%$. Specifically, for the variable extraction task, PromFuzz achieves a precision of $96.23\%$, with two false positive. The false positives are due to unconventional variable naming. For the statement extraction task, it achieves a precision of $83.33\%$, with only one false positive. This false positive arises from the complex business behaviors in smart contracts, where PromFuzz fails to differentiate between multiple similar behaviors in a single complex case. These critical variables and principal statements are expected to derive $26$ correct invariant checkers but ultimately generate $24$ correct ones, resulting in a precision of $92.31\%$.
TABLE IV: Variables and statements extraction accuracy.
@@
Variables Extraction 51 2 96.23%
Statements Extraction 5 1 83.33%
Overall 56 3 94.92%
IV-D Functional Bug Detection Accuracy
@@
This subsection answers RQ3. In this step, we evaluate the functional bug detection accuracy of PromFuzz with two metrics: recall and F1-score. We compare PromFuzz with four off-the-shelf tools: GPTScan, SMARTINV, ItyFuzz, and SMARTIAN [10]. We also consider a comparison between PromFuzz and PropertyGPT [22]. However, due to the unavailability of its code, we are unable to conduct this comparison. We do not compare PromFuzz with several program verification-based techniques [18, 28, 11], since they are largely ineffective in detecting functional bugs, as their constraints are designed for implementation bugs. Both ItyFuzz and SMARTIAN are fuzzers designed for detecting bugs in smart contracts. We perform them on $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$.
TABLE V: Functional bug detection accuracy.
Tool Overall
TP FP FN Recall F1-Score
PromFuzz 20 0 3 86.96% 93.02%
GPTScan 7 12 16 30.43% 33.33%
SMARTINV 6 43 17 26.09% 16.67%
ItyFuzz 0 0 23 0 0
SMARTIAN 3 0 16 15.79% 27.27%
@@
As shown in Table V, PromFuzz demonstrates superior performance in detecting functional bugs across three datasets. PromFuzz achieves an overall recall of 86.96% and an F1-score of 93.02%, detecting 20 true positives and maintaining 0 false positives across all datasets, with only 3 false negatives. Further analysis of the false negatives identified by PromFuzz reveals two primary causes. First, one false negative results from inaccuracies in the extraction of critical variables and principal statements. Second, the remaining two false negatives occur during the LLM-driven analysis step, where GPT fails to identify two critical bugs. Besides, with each smart contract project scan, PromFuzz takes an average of 43.8 seconds and costs approximately \$0.15, demonstrating its great efficiency and cost-effectiveness.
@@
In comparison, GPTScan and SMARTINV yield similar numbers of true positives and false negatives. However, SMARTINV records the highest number of false positives among the five tools assessed. Furthermore, ItyFuzz fails to identify functional bugs without the specific invariants we provide. SMARTIAN, another fuzzing tool, shows inconsistent performance as it fails to operate on several smart contract files. Consequently, we exclude SMARTIAN from testing against all bugs. We further analyze the false positives and false negatives of these tools. . For GPTScan, the primary reason arises from its single-agent design, which can lead to the incorrect identification of correct behaviors as functional bugs or the failure to detect genuine flaws. For SMARTINV, the primary reason arises from its lack of comprehensive reasoning capabilities, resulting in some bugs being overlooked or misidentified. Both ItyFuzz and SMARTIAN are limited by their original oracles, which lack the capability to identify functional bugs effectively.
IV-E Zero-day Functional Bug Detection
This subsection answers RQ4. We perform PromFuzz on $D_{Real-DeFi}$ to evaluate its performance in detecting zero-day functional bugs in real-world smart contracts. As illustrated in Table VI, PromFuzz successfully identifies $30$ zero-day functional bugs across $6$ of the $10$ Defi project. We have reported all detected bugs to the respective vendors, and $24$ of these have been assigned CVE IDs. The total market value of DeFi projects affected by these bugs is estimated at $\mathdollar 18.2$ billion. As these bugs have not yet been fully resolved, we maintain confidentiality by anonymizing the details of both the vendors and the specific CVE information. Moreover, we notice that zero-day bugs primarily fall into four categories: 13 bugs involve AMM price oracle manipulation, $7$ concern non-AMM price oracle manipulation, $4$ are related to wrong checkpoint order, $4$ pertain to risky first deposit, and $2$ involve wrong interest rate order. Further analysis in Section V delves into the practical impacts of these bugs, examining specific cases to illustrate how attackers exploit these bugs.
@@ IV-E Zero-day Functional Bug Detection
TABLE VI: Zero-day functional bug detection results. TVL represents the total value of cryptocurrency assets locked in the respective DeFi projects.
@@ 5 Yield & Sericves \$ 3,700M 5 5
V Case Study
To obtain an in-depth understanding of the practical impacts of functional bugs, we detail two such bugs identified by PromFuzz in real-world DeFi projects.
@@ V Case Study
5 if (_amount > type(uint128).max $|$ _amount == 0 $|$ share == 0){ revert InvalidAmount(); }
@@
10 if (totalPooledEther == 0){ return _depositAmount; }
@@
Figure 7: The Risky First Deposit bug (line 10).
Risky First Deposit. As shown in Figure 7, the code contains a risky first deposit bug that can lead to an unfair advantage for the first user and a distorted distribution of shares. Specifically, at line 10, when totalPooledEther == 0, the share variable at line 4 is set equal to _depositAmount, since the return value is assigned to _depositAmount in the _sharesForDepositAmount() function. Subsequent depositors have their shares calculated by (_depositAmount $\times$ eETH.totalShares()) / totalPooledEther, which depends on the values of eETH.totalShares() and totalPooledEther. While the calculation logic seems standard for compliant depositors, the first depositor can exploit the system by manipulating totalPooledEther. We can front-run other depositors’ transactions and inflate the price of pool tokens through a substantial “donation.” For instance, a malicious early user could deposit 1 wei of asset tokens as the first depositor, receiving 1 wei of shares tokens. Subsequently, the malicious sends 1 ETH to the pool, inflating totalPooledEther to 1 ETH + 1 wei (equal to ($1e18+1$) wei). In this scenario, the second depositor, depositing 2 ETH (equal to $2e18$ wei) of asset tokens, would receive only 1 wei of shares tokens. Because 1.99 wei of shares tokens (calculated by $2e18/(1e18+1)$) will be rounded down to 1 wei of shares tokens in smart contracts. They will lose 0.5 ETH if they withdraw right after the _deposit(). This disproportionate calculation of the first user’s shares can disadvantage subsequent users who receive fewer shares relative to their deposits.
AMM Price Oracle Manipulation. As shown in Figure 8, the code contains an AMM price oracle manipulation bug, which arises from the method of calculating token prices based solely on the supplies of two tokens. Specifically, _reserveA and _reserveB represent the respective reserve amounts of two assets. In line 13 of the code, the amount of token B is calculated using the amount of token A, along with the reserves of token A and token B. However, if the reserve amount of either token is maliciously manipulated, it can distort the price and lead to significant price fluctuations. For instance, consider token exchanges in the smart contract using the PoolHelpers library. Initially, the pool contains 500 token A and 500 token B. An attacker exchanges 100 token A for 100 token B, changing the pool to 600 token A and 400 token B. Then, the second user exchanges 30 token A and can only get 20 token B in return, leaving the pool with 630 token A and 380 token B. Finally, the attacker uses the 100 token B they obtained to exchange for 165 token A, ultimately making a profit of 65 token A.
@@
Figure 8: The AMM Price Oracle Manipulation bug (line 13).
VI Discussion
@@ VII Related work
Recent efforts to evaluate the security of smart contracts have produced various methods, categorized into static analysis, dynamic analysis, and AI-based analysis.
@@
Static Analysis. Tools like Slither [13], Zeus [18], Securify [36], Madmax [14], Invcon [21], Verismart [32], and Verisol [39] analyze smart contract code statically. Slither converts Solidity code into SlithIR to find bugs. Securify uses compliance and violation patterns to assess code safety. Madmax decompiles EVM bytecode and uses a logic-based specification for bug detection. Verisol translates Solidity into Boogie, formalizing semantic conformance against state machine workflows and reducing semantic checking to safety verification. However, these tools primarily focus on implementation bugs like integer overflow and reentrancy, and are less effective at identifying functional bugs.
@@
Dynamic Analysis. Tools like ItyFuzz [30], Smartian [10], Harvey [43], Echidna [15], ContractFuzzer [17], Mythril [5], Sailfish [8], Manticore [26], Teether [19], and Oyente [25] analyzes programs during execution. ItyFuzz uses snapshot-based fuzzing to reduce re-execution overhead, prioritizing intriguing snapshot states. Smartian combines static analysis for seed generation with feedback mechanisms. Sailfish identifies state-inconsistency bugs using a combination of lightweight exploration and symbolic evaluation. While dynamic analysis captures more realistic behaviors by executing smart contracts, it still primarily targets implementation bugs and struggles to effectively detect functional bugs.
@@
AI-based Analysis. SmarTest [31], ILF [16],xFuzz [44], and GNN-based tools [47, 23, 24] use AI to enhance smart contract bug detection. SmarTest combines symbolic execution with a language model to prioritize bugs. GNN-based tools leverage graph neural networks with expert patterns. While these tools improve detection capabilities, they mainly address implementation bugs. Recently, LLM-based tools like GPTScan [33] and SMARTINV [37] have been used to address functional bugs by analyzing discrepancies between business logic and code. These works, including the PromFuzz introduced here, open new avenues for detecting functional bugs in smart contracts using AI’s cognitive capabilities.
VIII Conclusion
@@
In this paper, we introduce PromFuzz, an automated and practical system designed to detect functional bugs in smart contracts, which bridges the gap between understanding the high-level business model logic and scrutinizing the low-level implementations in smart contracts. PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, representing a significant performance improvement over state-of-the-art methods. Additionally, we have created the first ground-truth dataset for functional bugs including $261$ contracts from various DeFi projects. Furthermore, our in-depth analysis of $10$ real-world DeFi projects has identified $30$ zero-day bugs, with $24$ of them receiving CVE IDs. Our discoveries have safeguarded assets totaling $\mathdollar 18.2$ billion from potential monetary losses.
Acknowledgment
This work was partly supported by the NSFC under No. U244120033, U24A20336, 62172243, 62402425 and 62402418, the Open Research Fund of Engineering Research Center of Blockchain Application, Supervision And Management (Southeast University), Ministry of Education, the China Postdoctoral Science Foundation under No. 2024M762829, the Zhejiang Provincial Natural Science Foundation under No. LD24F020002, the “Pioneer” and “Leading Goose” R&D Program of Zhejiang under 2025C01082 and 2025C02263, and the Zhejiang Provincial Priority-Funded Postdoctoral Research Project under No. ZJ2024001.